According to table 1.3, what volume of iron would be needed to balance a 1.00cm3 sample of lead on a two pan laboratory balance? Table 1.3 shows densities g/cm3 such as iron=7.87g/cm3 and lead is 11.4g/cm3
Answer #1
We have access to no textbook and therefore can not see table 1.3.

Sra
answered by: lilman
Answer #2
what time and financial constraints have you face since starting college? How did you deal with them
answered by: uma
Answer #3
density= m/v
7.87 g/cm3 = m/ 1.00cm3
7.87g / 11.36 g/cm3
answer is 0.69 cm3
answered by: fithgrader
Answer #4
Mass/Density = Volume

_11.4_g_ = v
7.87g/cm3

1.45cm3 = v
answered by: Nisha